OpenStudy (anonymous):

Evaluate the polynomial 6x - 7 for 1/2 n2-1/2n and y=2

OpenStudy (anonymous):

sorry, I cannot understand the equation.

OpenStudy (anonymous):

can u type it using the equation tab?

OpenStudy (anonymous):

there can i and what is that lol

OpenStudy (anonymous):

\[\frac{1}{2}n ^{2}-\frac{1}{2}n\] is it meant to be this?

OpenStudy (anonymous):

yes and y=2

OpenStudy (anonymous):

how did u type it like that?

OpenStudy (anonymous):

ok, the question seems weird I'm not sure if u edited it or not but it is phrased weirdly.

OpenStudy (anonymous):

Using the equation tab below, with the fraction part using \frac{xx}{yy}

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

do u hv a little time on ur hands i can show u the questin if u have time?

OpenStudy (anonymous):

sorry, I cannot help u with this question, it does not make sense to me. k sure, let me see it.

OpenStudy (anonymous):

I may be able to help u, but let me see the question first.

OpenStudy (anonymous):

okay, I think u should delete that right now, u are not suppose to do that.

OpenStudy (anonymous):

and what if other people see it?

OpenStudy (anonymous):

tell ur account and password.

OpenStudy (anonymous):

well do u have itif so i will

OpenStudy (anonymous):

yes I do, but please delete it now and that was not a smart move.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

u could have told me in msg, and I don't meant to be mean but please think before doing things next time.

OpenStudy (anonymous):

i did not think of that. and i will